OpenStudy (anonymous):

how to solve for x ? 5^(x-2) = 3^(3x+2)

OpenStudy (anonymous):

\[5^{x-2} =3^{3x+2}\]

OpenStudy (anonymous):

5^(x-2) = 3^(3x+2) => 5^x / 5^2 = (3^3)^x * 3^2 => 5^x = 225 * 27^x => (5/27)^x = 225 => x ln(5/27) = ln(225) => x = ln(225) / ln(5/27) => x ≈ - 3.21.

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

why ln instead of log ? @GirlByte

OpenStudy (anonymous):

In my opinion, I think this technique is better

OpenStudy (kira_yamato):

Actually it doesn't make a diff. because \[\frac{\log x}{\log y} = \frac{\left( \frac{\ln x}{\ln 10} \right)}{\left( \frac{\ln y}{\ln 10} \right)} = \frac{\left( \frac{\ln x}{\cancel{\color{red}{\ln 10}}} \right)}{\left( \frac{\ln y}{\cancel{\color{red}{\ln 10}}} \right)} = \frac{\ln x}{\ln y}\]